We collected some info on spectral peaks in the recent sensitivity. We used some results of Bruco run on the 45Mpc data of Dec. 6th https://scientists.virgo-gw.eu/DataAnalysis/bruco/2023/2023-12-06/Hrec_hoft_1385937603/bruco.html
(thanks to Ilaria and Maddalena who triggered this study)
- 113,114,120,123 Hz, 158, 163, 165 Hz (rounded to the nearest integer): have coherence with SDB1 controls, in particular channels SDB1_LC_TY (according to Michal Was, this signal is presently used as error signal for the SDB1 TY control actuating on the marionette) and also with accelerometers on the SDB1 vacuum chamber. Figure 1. All of them were seen also in the past (in particular during the lock of Oct. 27, purple curves in Figure 1) but after December 6th they are on average more intense (apart of the peak at 158 Hz which has always been present... it might not be the same peak though). The spectrogram in Figure 2 shows this.
- 184 Hz: this peak looks new since Dec 6th. As show in Figure 1, this peak has good coherence with the accelerometer on the SDB1 chamber. Note that a peak at this frequency got (very) excited in hrec during the tapping tests on SDB1 (Est flange and B5 viewport)
- 197 Hz: this peak is also more excited now than before Dec 6th, but has alwais been there. No coherence to report.
- 282 Hz: this peak looks new, and is present in Hrec since December 1st. It is coherent with seismic sensors. Some good ones are: the guralp on ceb floor (betwee BS and SR), accelerometers on SR and SPR: see Figure 3
- 309 Hz: this peak is NOT new (it was seen for example in the lock of October 27) but its maximum amplitude increased up to a factor 10 with respect to before. No coherence is reported. Figure 4.
- 370 Hz (is wide). More or less as the 309 Hz. They seem to move together. Figure 5.